Find the following:
a. the 11th term in the geometric sequence given that a1 = -3 and a2 = 9
b. the sum of the first 10 terms of the geometric sequence 5, 15, 45, ....
Thank you!

Answer by stanbon(75887)

Find the following:
a. the 11th term in the geometric sequence given that a1 = -3 and a2 = 9
-------
r = 9/(-3) = -3
----
A(11) = a(1)*r^(11-1) = -3*(-3)^10 = (-3)^11 = -177147
------------------
b. the sum of the first 10 terms of the geometric sequence 5, 15, 45, ....
a(1) = 5
r = 15/5 = 3
------
a(10) = 5*3^9 = 98415
----
S(n) = a(1)[(r^n - 1)/(r-1)
S(10) = 5[98414/2] = 5*49207 = 246035
-----
Cheers,
Stan H.
------------